pose Megan is an avid reader and buys only comic books. Megan deposits $4,000 in a bank account that pays an annual nominal interest rate of 10%. Assume this interest rate is fixed—that is, it won't change over time. At the time of her deposit, a comic book is priced at $20.00. Initially, the purchasing power of Megan's $4,000 deposit is 200 comic books. For each of the annual inflation rates given in the following table, first determine the new price of a comic book, assuming it rises at the rate of inflation. Then enter the corresponding purchasing power of Megan's deposit after one year in the first row of the table for each inflation rate. Finally, enter the value for the real interest rate at each of the given inflation rates. Hint: Round your answers in the first row down to the nearest comic book. For example, if you find that the deposit will cover 20.7 comic books, you would round the purchasing power down to 20 comic books under the assumption that Megan will not buy seven-tenths of a comic book. Annual Inflation Rate 0% 10% 13% Number of Comics Megan Can Purchase after One Year Real Interest Rate % % % When the rate of inflation is less than the interest rate on Megan's deposit, the purchasing power of her deposit over the course of the year.
Answer #1

Deposit= $4000

Nominal interest rate=10%

Price of comic books= $20.

Purchasing power of Meghan's deposit is 4000/20= 200 comic books.

Inflation rate→ 0% 10% 13%
Number of comics 4400/20= 220 4400/22=200 4400/22.6≈ 195
Real interest rate 10%-0%=10% 10%-10%=0% 10%-13%=-3%

Real interest rate= Nominal interest rate- Inflation

With 0% inflation rate

Deposit= 4000*(10%+0)= $4400

Price of comic book= $20

With 10% inflation rate

Deposit= 4000*1.1= $4400

Price of comic book= $20*1.1= $22

With 13% inflation rate

Deposit= 4000*1.1= $4420

Price of comic books=$20*1.13=$ 22.6

When the rate of inflation is less than the interest rate on Megan's deposit, the purchasing power of her deposit RISES over the course of the year.

If it helps kindly upvote